Supplemental Operations Report – March 5th

Supplemental Operations Report 31-Mar-2023

Name of person filing report: Shannon Rupert

Reason for Report: Changing of the guard aka site personnel. Tomorrow Sergii will be The Last Man on Mars.

Non-Nominal Systems: Director’s trailer toilet water pump is leaking
Action taken for non-nominal systems: Ordered new part and will turn off trailer water pump when I leave

Generator: Generator was serviced this morning
Solar–SOC Last 24 hours: We had rain and clouds most of the day yesterday but today was sunny.
Max 100% Min 41 % Avg 61.6%
Power System Notes: Generator schedule shortened to account for greater isolation in spring. Off around 7, on around 10, and this could be even shorter most days.

Propane Reading Station Tank: Filled today
Propane Reading Director Tank: Filled today
Propane Reading Intern Tank: Filled today
Propane Reading Generator Tank: Filled today

Water (Static Tank) – 250 gallons
Water in GreenHab – 50 gallons
Water (Outpost tank) – 480 gallons
Water in Science Dome: 0 gallons
Hab Toilet Tank emptied: n/a

Sojourner rover used: Yes
Hours: not checked
Beginning Charge: 100
Ending Charge: 41
Currently Charging: Yes
Notes on Rovers: All are plugged in, charged, and ready for service.

ATV’s Used: (Honda, 350.1, 350.2, 300): None
Reason for use: n/a
Oil Added? No
ATV Fuel Used: 0 Gals
Ethanol Free Gasoline — 0 Gallons
# Hours the ATVs were Used today: 0
Notes on ATV: There was no time for me to get to Richfield (nor good enough weather) to purchase the trailer. I will be doing that remotely and Sergii will pick it up.

HabCar used and why, where? Yes. To town and also by the crew, tonight it was hooked up to the trailer as tomorrow the dump is open, but the battery died so it is currently charging overnight. This is the second time the battery has died after it has been started and operating, so we don’t know what’s going on at this time.

Crew Car used and why, where? Yes. To town and also by the crew. This crew will be taking it back to GJ tomorrow so that the next crew will be able to pick it up.
Luna Used and why, where? I will be taking Luna back to WA with me tomorrow.

General notes and comments: Crew 276 has set up their new “Thunderdome” in the same area as last year. We moved the student camp from the previous spot south and east of the station to a more protected site north of the Hab, just off Hab Ridge on the flat. Things have gone smoothly.

Dual Split in Science Dome: off due to warmer weather.

Summary of Internet: Nominal

Summary of suits and radios: nothing to report

EVA COMMS: Used by faculty to communicate with each other and also with on site personnel

Campus wide inspection, if action taken, what and why: nothing to report

Summary of General Operations: Crew 276 on site has about 27 faculty and students in simi-sim. Correction from last week—the rovers were not serviced as reported.

Summary of Hab operations: Nominal

Summary of GreenHab Operations: I completed the overhaul of the GreenHab today. About 1/3 of the tomatoes were removed and the remaining are still producing. After this next crew, the tomatoes in the herb garden should be removed. I removed the beans as they were no longer producing. I set up the red wagon to remix soil for the end of the season.

On Tuesday, I harvested 593 g of tomatoes. Tonight I asked any of the crew who were interested to go harvest what was ripe, an estimate of another 500 g.

I lowered the thermostat to 55 and it is fine to be set there for the remainder of the season. I also suggest the door stay open on days where it is above 55 and the wind is not high.

There is a rodent in one of the greens beds. It seems to like beet greens. We will be setting a life trap for this little beet fiend tomorrow. It may be the same visitor that has been running over people in the Hab.

Summary of SciDome Operations: The leaks from the windowsill are getting worse and the black mold on the inside of the dome continues to grow. We will be looking at it as soon as things dry out from the constant battle between winter and spring as to what season it actually is.

Summary of any Observatory Issues: Photos taken of the twisted belt and sent to the Robotic Observatory dome company

Summary of RAM Operations: Nothing to report

Summary of Outpost Operations: I moved out of my trailer today. Most traces of my existence here have been packed and loaded into the van (except my rocks—I’m going to need to pick them up in May). Mostly empty now, the trailer reminds me of the day I first saw it and knew instantly it was the one to be my home. I paid for it with money my grandmother had left me, and I’ve donated it to TMS. My grandmother would have been confused, but supportive. It’s been a safe and cozy little home for me here. Both trailers also have new tabs.

Summary of Health and Safety Issues: Nothing to report

Questions, Concerns, Supplies needed and Requests: Nothing to report

Copyright © The Mars Society. All rights reserved. | Main Site